# Enough Coin Flips Can Make LLMs Act Bayesian

### v1

@@ Abstract
Large language models (LLMs) exhibit the ability to generalize given few-shot examples in their input prompt, an emergent capability known as in-context learning (ICL). We investigate whether LLMs utilize ICL to perform structured reasoning in ways that are consistent with a Bayesian framework or rely on pattern matching. Using a controlled setting of biased coin flips, we find that: (1) LLMs often possess biased priors, causing initial divergence in zero-shot settings, (2) in-context evidence outweighs explicit bias instructions, (3) LLMs broadly follow Bayesian posterior updates, with deviations primarily due to miscalibrated priors rather than flawed updates, and (4) attention magnitude has negligible effect on Bayesian inference. With sufficient demonstrations of biased coin flips via ICL, LLMs update their priors in a Bayesian manner.
@@ 1 Introduction
Large language models (LLMs) designed for next-token prediction have gained significant popularity, largely because of their ability to generalize beyond language prediction and perform a wide range of novel tasks without requiring explicit weight updates Brown et al. (2020). Methods to induce emergence in controlled ways include techniques such as chain-of-thought prompting Wei et al. (2022), prompt chaining Wu et al. (2022), and in-context learning (ICL). ICL, particularly, provides demonstrations of a specific task to the model as part of its input prompt.
@@ 2 Background & Related Work
Brown et al. (2020) introduces in-context learning (ICL) as a mechanism for few-shot generalization in language models. Although ICL usage has surged, users rarely employ it as a method to align models with target distributions. Further, issues with models’ sensitivity to the positioning of tokens in their prompts have complicated the effective use of ICL as an alignment technique. Lu et al. (2022) demonstrates that the positioning of information within an ICL prompt affects model performance and devises a permutation-based approach to overcome this bias. Liu et al. (2023) extends this analysis to highlight a persistent “lost-in-the-middle” effect, in which the information in the middle of a prompt is down-weighted.
Our work shows that in-context rollouts of a probability distribution correlate well with the mean of a Bayesian posterior, and we further show that LLMs have a time-varying discount factor over the ICL prompt.
@@
Many authors have explored the mechanism by which ICL emerges in language models. Xie et al. (2021) finds that ICL can be viewed as a language model implicitly performing Bayesian inference—i.e., ICL emerges via modeling long-range coherence during pretraining. Jiang (2023) shows that emergent capabilities of LLMs, such as ICL, are Bayesian inference on the sparse joint distribution of languages. Wang et al. (2024) react to the ordering sensitivity of ICL prompts and pose ICL as a natural side effect of LLMs functioning as latent variable models. Finally, Zhang et al. (2023) posit that ICL is an implicit form of Bayesian model averaging.
@@
Our own findings, that LLMs can often apply Bayesian-like updates despite relying on miscalibrated priors, resonate with this need for calibration, underscoring the importance of correcting initial biases when using LLMs in downstream tasks. We confirm the ordering sensitivity of ICL prompts and further show empirically that ICL has several implicit Bayesian modeling behaviors. Finally, we demonstrate that it is unlikely that attention magnitude is a key component of the formalization.
@@ 4 Understanding the LLM Prior
Scaling the language model size has shown effectiveness in many tasks. Therefore, we explore whether scaling also boosts performance on modeling expected biased distribution. We use Pythia Scaling Suite Biderman et al. (2023) that covers model size ranging from 70M to 12B and test on different biased $\theta$. Results from this experiment are presented in Figure 4. For a given bias, scaling the model size does not substantially change the language models’ priors or improve the performance of modeling expected distributions. However, the relative ordering among different biases does shift as the model size increases.
@@ 5.3 Does attention impact updates?
Some prior work, such as Zhang et al. (2023), suggests that attention helps to weight the Bayesian update. In this section, we aim to leverage our simplified setup to empirically understand the impact that attention has on the convergence behavior of the model. We use the same setup as subsection 5.2 but instead draw the center $K$ samples $\sim Binom(K,\theta_{1})$ and the outer $M=100-K$ samples $\sim Binom(M,\theta_{2})$.
Figure 8 plots the relationship between total attention and model point-estimate extremity under the Bayesian posterior ($\gamma=1.0$) (i.e. the value of the CDF of the true posterior at the model point estimate). We can see that the amount of attention paid to any segment is generally uncorrelated with the overall quality of the point estimate ($\theta_{1}:(R=0.02,p=0.48),\>\theta_{2}:(R=-0.03,p=0.36)$), suggesting that the total magnitude of the attention paid to each segment does not dramatically impact model quality.
Figure 8: Relationship between total attention and model point-estimate extremity under the Bayesian posterior ($\gamma=1.0$). Overall, the extremity of the model point estimate under the Bayesian model appears uncorrelated with the attention.
In addition, the fraction of attention has a similar lack of correlation, as shown in Figure 9, which suggests that paying any special attention (in terms of magnitude) to any particular ICL example is uncorrelated with downstream performance during model updates.
Figure 9: Fraction of attention assigned to samples from $\theta_{1}$ versus the deviation between the model-predicted distribution and the true posterior mean for Llama-3.1-8B. The findings suggest that the relative attention paid to in-context examples does not directly predict the model’s update performance.
@@ 7 Limitations
Another methodological limitation arises in evaluating closed-source models. Since our approach relies on extracting logits to approximate posterior distributions, it cannot be directly applied to black-box models such as GPT-4 or Claude, where such internals are inaccessible. While an alternative approach using sampling via API calls could approximate the posterior, this method is costly and susceptible to distortions from API-side interventions such as caching, response smoothing, or temperature adjustments. These factors could introduce artifacts that obscure the model’s true Bayesian reasoning capabilities.
@@ Acknowledgments
As part of their affiliation with UC Berkeley, the authors were supported in part by the National Science Foundation, the U.S. Department of Defense, the Ford Foundation, and/or the Berkeley Artificial Intelligence Research (BAIR) Industrial Alliance program. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of any supporting entity, including the Department of Defense or the U.S. Government.
@@
  • Appendix F explores similar results to section 4 and section 5 but with dice rolling (as opposed to coin flips).
@@ E.2 ICL Scaling Results
Here we present all the results from the ICL scaling experiments in Section 5.1.
@@
Appendix F Rolling Dice
@@
The results are shown in Figure F.8. For each bias percentage, we averaged results across the six die faces and 50 prompt variants, totaling 300 trials per bias percentage. Non-instruct models generally performed better than their instruct counterparts, and best around a 50%-60% bias, struggling more with higher biases. Instruct model performance was more varied, with some models showing little change in behavior and others improving as the bias value increased.
@@
In Figure F.3, Figure F.4, Figure F.5, Figure F.6, and Figure F.7, we present ICL plots measuring TVD for a variety of model variants on the simple dice rolling experiment. These results correlate well with the results observed in section 4, the coin flip experiments.
Figure F.1: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Llama3.1-8B model.
Figure F.2: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Llama3.1-8B-Instruct model.
Figure F.3: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Microsoft Phi-2 model.
Figure F.4: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Microsoft Phi-3.5-mini-instruct model.
Figure F.5: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Google Gemma-2-2B model.
Figure F.6: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Mistral-7B-Instruct model.
Figure F.7: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Mistral-7B model.
Figure F.8: Biased die rolls: Plots of mean total variation distance (TVD, $\downarrow$) against bias percentage for non-instruct (left) and instruct (right) models when aggregated across prompts (N=50) for the biased die rolling experiment.

### v2

@@ Abstract
Large language models (LLMs) exhibit the ability to generalize given few-shot examples in their input prompt, an emergent capability known as in-context learning (ICL). We investigate whether LLMs use ICL to perform structured reasoning in ways that are consistent with a Bayesian framework or rely on pattern matching. Using a controlled setting of biased coin flips, we find that: (1) LLMs often possess biased priors, causing initial divergence in zero-shot settings, (2) in-context evidence outweighs explicit bias instructions, (3) LLMs broadly follow Bayesian posterior updates, with deviations primarily due to miscalibrated priors rather than flawed updates, and (4) attention magnitude has negligible effect on Bayesian inference. With sufficient demonstrations of biased coin flips via ICL, LLMs update their priors in a Bayesian manner. Code and visualizations are available on the project page.
@@ 1 Introduction
Large language models (LLMs) designed for next-token prediction have gained significant popularity, largely because of their ability to generalize beyond language prediction, and perform a wide range of novel tasks without requiring explicit weight updates Brown et al. (2020). Methods to induce emergence in controlled ways include techniques such as chain-of-thought prompting Wei et al. (2022), prompt chaining Wu et al. (2022), and in-context learning (ICL). ICL, particularly, provides demonstrations of a specific task to the model as part of its input prompt.
@@ 2 Background & Related Work
Brown et al. (2020) introduces in-context learning (ICL) as a mechanism for few-shot generalization in language models. Although ICL usage has surged, users rarely employ it as a method to align models with target distributions. Further, issues with models’ sensitivity to the positioning of tokens in their prompts have complicated the effective use of ICL as an alignment technique. Lu et al. (2022) demonstrates that the positioning of information within an ICL prompt affects model performance and devises a permutation-based approach to overcome this bias. Liu et al. (2023) extends this analysis to highlight a persistent “lost-in-the-middle” effect, in which there is implicit positional bias for information as it relates to accuracy and suggest that the mechanism behind the lost-in-the-middle effect may be more closely related to position embedding. Liu et al. (2024) show that LLMs can extrapolate the behavior of dynamical systems given large numbers of in-context examples. However, their discovered power-law fits imperfectly, demonstrating high loss at long contexts.
Our work explores a time-varying discount factor for in-context learning, more directly explaining the higher-than-expected loss at long context lengths. We demonstrate that in-context rollouts of a probability distribution correlate well with the mean of a Bayesian posterior. Further, we analyze how attention weights affect output accuracies and find little correlation.
@@
Many authors have explored the mechanisms through which ICL capability emerges in language models. Xie et al. (2021) finds that ICL can be viewed as a language model implicitly performing Bayesian inference—i.e., ICL emerges via modeling long-range coherence during pretraining. Jiang (2023) shows that emergent capabilities of LLMs, such as ICL, are Bayesian inference on the sparse joint distribution of languages. Wang et al. (2024) react to the ordering sensitivity of ICL prompts and pose ICL as a natural side effect of LLMs functioning as latent variable models. Finally, Zhang et al. (2023) posit that ICL is an implicit form of Bayesian model averaging.
@@
Our own findings, that LLMs can often apply Bayesian-like updates despite relying on miscalibrated priors, resonate with this need for calibration, underscoring the importance of correcting initial biases when using LLMs in downstream tasks. We confirm the ordering sensitivity of ICL prompts and further show empirically that ICL has several implicit Bayesian modeling behaviors. Finally, we demonstrate that it is unlikely that attention magnitude is a key component of this formalization.
@@ 4 Understanding the LLM Prior
Scaling the language model size has shown effectiveness in many tasks. Therefore, we explore whether scaling also boosts performance on modeling expected biased distribution. We use the Pythia Scaling Suite Biderman et al. (2023), which covers model sizes ranging from 70M to 12B, and test on different biased $\theta$. Results from this experiment are presented in Figure 4. For a given bias, scaling the model size does not substantially change the language models’ priors or improve the performance of modeling expected distributions. However, the relative ordering among different biases does shift as the model size increases.
@@ 5.3 Does attention impact updates?
Some prior work, such as Zhang et al. (2023), suggests that attention helps to weight the Bayesian update. In this section, we aim to leverage our simplified setup to empirically understand the impact that attention has on the convergence behavior of the model. We use the same setup as Section 5.2 with a sequence $L$ of length $N=100$. There is a “switchover” point $K=50$ such that samples $L_{1-K}\sim\operatorname{Binom}(K,\theta_{1})$ and $L_{K-N}\sim\operatorname{Binom}(N-K,\theta_{2})$. We experiment varying $K\in[10,90]$.
Figure 8 plots the relationship between total attention and model point-estimate extremity under the Bayesian posterior ($\gamma=1.0$) (i.e. the value of the CDF of the true posterior at the model point estimate) for all $K$. We can see that the amount of attention paid to any segment is generally uncorrelated with the overall quality of the point estimate ($\theta_{1}:(R=0.02,p=0.48),\>\theta_{2}:(R=-0.03,p=0.36)$), suggesting that the total magnitude of the attention paid to each segment does not dramatically impact model quality.
Figure 8: Relationship between total attention and model point-estimate extremity under the Bayesian posterior ($\gamma=1.0$), and all values of $K$. Overall, the extremity of the model point estimate under the Bayesian model appears uncorrelated with the attention.
In addition, the fraction of attention, for all $K$, has a similar lack of correlation, as shown in Figure 9, which suggests that paying any special attention (in terms of magnitude) to any particular ICL example is uncorrelated with downstream performance during model updates. Additionally, there is no significant difference in results as we vary $K$, visualized in Appendix F.
Figure 9: Fraction of attention assigned to samples from $\theta_{1}$ versus the deviation between the model-predicted distribution and the true posterior mean for Llama-3.1-8B for all values of $K$. The findings suggest that the relative attention paid to in-context examples does not directly predict the model’s update performance.
@@ 6 Discussion & Conclusion
However, it is worth asking if LLM probabilities ought to be calibrated at all? While we primarily focus on the mechanism in this work, i.e., adjusting LLM probabilities with in-context evidence, we believe that LLMs used as agents should be well-calibrated. One of the primary reasons for this is the growing adoption of LLMs in simulation, particularly probabilistic simulations and world modeling, in which it is quite important to correctly model stochastic outcomes. In addition, well-calibrated models will likely make more fair/unbiased decisions than uncalibrated models Tian et al. (2023).
In future work, we would like to explore how our work’s discoveries map to multimodal language models. Prior work has shown that vision-language models (VLMs) are blind Rahmanzadehgervi et al. (2025) and fail to perform on tasks that are dominated by simple visual reasoning. Petryk et al. (2024) attempted to measure this misalignment in VLMs by analyzing hallucinations in image captioning. Evidently, VLMs fail to accurately correlate visual features with textual prompts, pointing towards hidden miscalibration. Exploring purely visual stochastic tasks and how VLMs perform in those settings is a natural extension to this work.
@@ 7 Limitations
Another methodological limitation arises in evaluating closed-source models. Since our approach relies on extracting logits to approximate posterior distributions, it cannot be directly applied to black-box models such as GPT-4 or Claude. While an alternative approach using sampling could approximate the posterior, this method is costly and susceptible to distortions from API-side interventions such as caching, response smoothing, or temperature adjustments introducing artifacts that obscure the model’s true Bayesian reasoning capabilities.
@@ Acknowledgments
This paper would not be possible without the Berkeley AI Research (BAIR) espresso machine. It was huddled around this machine did we come up with the idea for this line of questioning in the first place. We would also like to thank Anand Siththaranjan for their review of the paper and providing excellent feedback to make this work better.
As part of their affiliation with UC Berkeley, the authors were supported in part by the National Science Foundation, the Ford Foundation, and/or the Berkeley Artificial Intelligence Research (BAIR) Industrial Alliance program. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA), the Army Contracting Command-Aberdeen Proving Grounds (ACC-APG), and the Air Force Research Laboratory under Contract No(s) W912CG-24-C-0011, FA8650-23-C-7316. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of any supporting entity, including the AFRL, ACC-APG, the Department of Defense or the U.S. Government.
@@
  • Appendix F presents additional figures demonstrating the impact of varying $K$, the switchover point.
  • Appendix G explores similar results to section 4 and section 5 but with dice rolling (as opposed to coin flips).
@@ E.2 ICL Scaling Results
In Figure E.7, we present all the results from the ICL scaling experiments in Section 5.1.
@@
Appendix F Varying the Switchover Point
We also perform several experiments varying the value $K$ (the switchover point) in the experiments. Figure F.1 shows the correlation between the amount of attention paid within the cutoff region and the calibration accuracy, where we see that while the size of the cutoff ($K$) does impact the amount of attention paid to the model, there is little correlation between that amount of attention and the calibration accuracy. Similarly, Figure F.2 shows the correlation between the amount of attention paid outside the cutoff region and the calibration accuracy, demonstrating a similar lack of correlation.
These results are further shown in Figure F.3 which plots the deviation of the value $\theta$ against the expected Bayesian update probability for different values of $K$. We can see that as the probabilities become more extreme, the deviation becomes higher, and models have more trouble adjusting to more extreme probabilities, however there is no statistically significant difference between the $K$ values.
Figure F.1: This plot shows the correlation between the amount of attention paid within the cutoff region and the calibration accuracy.
Figure F.2: This plot shows the correlation between the amount of attention paid outside the cutoff region and the calibration accuracy.
Figure F.3: Plot showing the deviation of the model predicted $\theta$ against the expected Bayesian update probability for different values of $K$.
Appendix G Rolling Dice
@@
The results are shown in Figure G.8. For each bias percentage, we averaged results across the six die faces and 50 prompt variants, totaling 300 trials per bias percentage. Non-instruct models generally performed better than their instruct counterparts, and best around a 50%-60% bias, struggling more with higher biases. Instruct model performance was more varied, with some models showing little change in behavior and others improving as the bias value increased.
@@
In Figure G.3, Figure G.4, Figure G.5, Figure G.6, and Figure G.7, we present ICL plots measuring TVD for a variety of model variants on the simple dice rolling experiment. These results correlate well with the results observed in section 4, the coin flip experiments.
Figure G.1: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Llama3.1-8B model.
Figure G.2: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Llama3.1-8B-Instruct model.
Figure G.3: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Microsoft Phi-2 model.
Figure G.4: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Microsoft Phi-3.5-mini-instruct model.
Figure G.5: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Google Gemma-2-2B model.
Figure G.6: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Mistral-7B-Instruct model.
Figure G.7: Mean total variation distance (TVD, $\downarrow$) vs. bias percentage for several ICL example lengths on the die rolling task for the Mistral-7B model.
Figure G.8: Biased die rolls: Plots of mean total variation distance (TVD, $\downarrow$) against bias percentage for non-instruct (left) and instruct (right) models when aggregated across prompts (N=50) for the biased die rolling experiment.